# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v1

@@ B.1 Specific flood events
Code Location Date # homes in aff. mun. # homes at risk # hit homes
@@
EMSR260 North-West area 12th Dec. 2017 5825 4607 2
EMSR359 Emilia-Romagna 15th May 2019 5514 4185 1
EMSR483 Calabria 22nd Nov. 2020 379 94 0
EMSR486 Sardegna 28th Nov. 2020 121 8 0
EMSR487 Emilia-Romagna 6th Dec. 2020 2161 1833 4
EMSR488 Lazio 9th Dec. 2020 149 72 0
EMSR649 Sicilia 9th Feb. 2023 2534 159 1
@@
Flag multi floor $0.0^{***}$ $0.0^{***}$
@@
EMSR192 Logarithm of square meters $0.076^{*}$ $0.023^{**}$
@@
Flag garage $0.013^{**}$ $0.013^{**}$
Flag air conditioning $0.074^{*}$ $0.086^{*}$
Flag annex $0.0^{***}$ $0.0^{***}$
@@
Flag multi floor $0.01^{***}$ $0.0^{***}$
@@ B.2 Hedonic regressions
We performed some complementary analyses to validate the specifications chosen for our main regressions. We present four different sets of supplementary results: (i) spatial correlation analysis to check if spatial econometric approaches are needed; (ii) robustness analysis for the risk zone definition; (iii) robustness analysis for the home characteristics included in the baseline regression, (iv) robustness of the historical memory specification of regions.
@@
Table A4: Test for spatial autocorrelation on the residuals of our baseline regression (corresponding to Equation 1). Note that we estimated the Global Moran’s I at the municipality level assuming a queen contiguity relationship between municipalities (Anselin et al., 2009). Significance levels: *** = p-value$<0.01$, ** = p-value $<0.05$, * = p-value $<0.1$.
@@
Figure A8 shows the effect on the home price of being in a risk zone across Italy for all parameter configurations. Flood risk causes a statistically significant price reduction, ranging between 1% and 5%. Once we focus on frequently, moderately, and rarely flooded regions, the largest price reductions are observed for the frequently flooded regions. Instead, coefficients for the moderately and rarely flooded regions are generally smaller and less significant than those for frequently flooded regions. The yellow circle in the plot marks the specification selected for the main text, which considers: the logarithm of home size, binned floors, fixed effects at OMI microzone, and the outliers have been removed. We can observe that the results shown in the main are representative also of other parameter specifications since they are in the bulk of points. The exception is the chunk of significant negative and positive coefficients for moderately and rarely flooded regions. These coefficients are associated with the municipality fixed effects as confirmed in Figure A9, and so probably highlight other factors than flood risk.
Figure A8: Effect on home prices of being in a flood-risk area (identified using ISPRA risk maps, first-left box) across Italy and according to the disaggregation in frequently, moderately, and rarely flooded regions based on the flood events recorded by EM-DAT (2024), second to fourth boxes. Each dot represents a parameter configuration, the color reports the statistical significance. The dots with the yellow edges identify the parameter configuration presented in the main text. R2 [min, max] = [0.48, 0.85]; observations [min, max] = [552761, 555880].
@@
Construction year: 1955-1960 (ref.: $<$1955) 0.021 0.003 ***
Construction year: 1960-1965 (ref.: $<$1955) -0.000 0.004
Construction year: 1965-1970 (ref.: $<$1955) 0.045 0.003 ***
Construction year: 1970-1975 (ref.: $<$1955) 0.050 0.004 ***
Construction year: 1975-1985 (ref.: $<$1955) 0.084 0.004 ***
Construction year: 1985-1995 (ref.: $<$1955) 0.175 0.004 ***
Construction year: 1995-2005 (ref.: $<$1955) 0.270 0.004 ***
Construction year: 2005-2015 (ref.: $<$1955) 0.343 0.004 ***
Construction year: 2015-2025 (ref.: $<$1955) 0.404 0.006 ***
Construction year: Missing (ref.: $<$1955) 0.046 0.013 ***
@@
To validate our definition of the historical memory of regions we repeated the hedonic regression analyses in Table 2 after considering two other measures of flood frequency: (1) an alternative dataset for historic flood events compiled from ISPRA records; (2) considering the normalization of the number of events by region area.
EM-DAT is well-known to over-represent disasters in richer areas (Felbermayr and Gröschl, 2014). Therefore, we compiled a dataset of flood events from ISPRA data and re-estimated our hedonic regressions. ISPRA collects data on flood events drawing on multiple sources, such as the Italian Civil Protection Department, ISTAT, the National Research Council, local administrations and newspapers. For each event, it reports the duration, an estimation of the damages, an estimation of the allocated aid funds, and the total number of deaths and missing or evacuated people. Although data are provided back to 1951, systematic data collection only began in 2003, so we only consider the later period for better accuracy. Data are not collected in tabular format, but we manually compiled a database reading from the yearly reports by ISPRA. [^44] https://indicatoriambientali.isprambiente.it/it/edizioni-annuario. The latest available report is for the year 2022. The ISPRA dataset contains 402 event-region pairs distributed across regions and macroareas. Since the inclusion criteria for ISPRA are less strict than EM-DAT, this results on some variations in the historical memory classifications. Even with these differences we still observe the same pattern of significance of the risk coefficients disaggregated by regions that was observed in the regression presented in the main manuscript (Table A7).
Table A7: Risk effect on home prices. For each spatial fixed effects, Models 1-3-5 report the effect of the risk on home prices estimated using Equation (1), while Models 2-4-6 disaggregate the effect of flood risk by the historical memory of the regions according to ISPRA. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
Risk -0.029∗∗∗ -$0∗∗$008∗∗
(0.010) (0.003) (0.003)
Risk $\times$ Frequently flooded regions -0.032∗∗ -0.016∗∗∗ -0.012∗∗∗
(0.015) (0.005) (0.005)
Risk $\times$ Moderately flooded regions -0.038∗∗∗ -0.001 -0.003
(0.012) (0.005) (0.005)
Risk $\times$ Rarely flooded region 0.051∗∗ 0.006 0.000
(0.021) (0.010) (0.013)
Observations 552761 552761 552761
@@
Spatial fixed effects Municipality OMI zone Census Tract
Number of clusters 7203 20297 160630
@@
If one assumes flood risk is at least to some extent uniformly spread throughout the Italian territory, the most affected regions would tend to be the largest ones. Thus, we repeated the hedonic regression analyses in Tables 2 and A7 using the historical memory levels defined through the area normalization. As expected, significant price reductions are observed in frequently flooded regions (Table A8).
Table A8: Effect of flood risk disaggregated across historical memory regions. The table compares the results considering the historical memory classification for EM-DAT and ISPRA data. In this case, the number of flood events is normalized by the region surface area. ***: p-value$<$0.01, **: p-value$<$0.05, *: p-value$<$0.10.
Dep. var.: log(P)
Model id 1 2 3 4 5 6
Dataset EM-DAT ISPRA EM-DAT ISPRA EM-DAT ISPRA
Normalization
Risk $\times$ Frequently flooded regions -0.044∗∗ -0.006 -0.017∗∗∗ -0.019∗∗∗ -0.019∗∗∗ -0.017∗∗∗
(0.021) (0.014) (0.004) (0.005) (0.005) (0.006)
Risk $\times$ Moderately flooded regions -0.032∗∗∗ -0.049∗∗∗ -0.008 -0.008 -0.004 -0.008
(0.010) (0.014) (0.006) (0.005) (0.006) (0.005)
Risk $\times$ Rarely flooded regions 0.022 0.023 0.005 0.004 0.006 0.005
(0.012) (0.014) (0.005) (0.006) (0.008) (0.008)
Observations 552761 552761 552761
R2 0.669 0.755 0.848
Spatial fixed effects Municipality OMI zone Census Tract
Number of clusters 7203 20297 160630
Clustered s.e.

### v2

@@ A.1 A more detailed description
Young Buyers’ Mortgage (Mutuo Giovani)
No 329705 (59.65%)
Yes 223023 (40.35%)
Missing 0 (0%)
@@ B.1 Specific flood events
Code Location Date
# homes
in aff. mun.
# homes at
risk
# hit
homes
@@
EMSR260 North-West area 12th Dec. 2017 5825 4607 $\leq 4$
EMSR359 Emilia-Romagna 15th May 2019 5514 4185 $\leq 4$
EMSR483 Calabria 22nd Nov. 2020 379 94 $\leq 4$
EMSR486 Sardegna 28th Nov. 2020 121 8 $\leq 4$
EMSR487 Emilia-Romagna 6th Dec. 2020 2161 1833 $\leq 4$
EMSR488 Lazio 9th Dec. 2020 149 72 $\leq 4$
EMSR649 Sicilia 9th Feb. 2023 2534 159 $\leq 4$
@@
Flag multi floor $^{***} 0.00.0^{***}$
@@
EMSR192 Logarithm of square meters $076^{*} 0.0$$023^{**}$
@@
Flag garage $^{**} 0.0130.013^{**}$
Flag air conditioning $074^{*} 0.0$$086^{*}$
Flag annex $^{***} 0.00.0^{***}$
@@
Flag multi floor $1^{***} 0.00.0^{***}$
@@ B.2 Hedonic regressions
We performed some complementary analyses to validate the specifications chosen for our main regressions. We present four different sets of supplementary results: (i) spatial correlation analysis to check if spatial econometric approaches are needed; (ii) robustness analysis for the risk zone definition; (iii) robustness analysis for the home characteristics included in the baseline regression, (iv) robustness of the historical memory (awareness) specification.
@@
Table A4: Test for spatial autocorrelation on the residuals of our baseline regression (corresponding to Equation 2). Note that we estimated the Global Moran’s I at the municipality level assuming a queen contiguity relationship between municipalities (Anselin et al., 2009). Significance levels: *** = p-value<0.01<0.01, ** = p-value <0.05<0.05, * = p-value <0.1<0.1.
@@
Figure A8 shows the effect on the home price of being in a risk zone across Italy for all parameter configurations and half memory time (see Section 2.4) of 10 years. Flood risk causes a statistically significant price reduction, ranging between 1% and 5%. Once we focus on transaction with high, medium, and low awareness, the largest price reductions are observed for the high awareness purchases. Instead, coefficients for the medium and low awareness transactions are generally smaller and less significant than those for high awareness. The yellow circle in the plot marks the specification selected for the main text, which considers: the logarithm of home size, binned floors, fixed effects at OMI microzone, and the outliers have been removed. We can observe that the results shown in the main are representative also of other parameter specifications since they are in the bulk of points. The exception is the chunk of significant negative coefficients for medium and low awareness transactions. These coefficients are associated with the municipality fixed effects as confirmed in Figure A9, and so probably highlight other factors than flood risk.
Figure A8: Effect on home prices of being in a flood-risk area (identified using ISPRA risk maps, first-left box) across Italy and considering high, medium, and low awareness transactions based on the flood events recorded by EM-DAT (2024) and $\tau=10$ years, second to fourth boxes. Each dot represents a parameter configuration, the color reports the statistical significance. The dots with the yellow edges identify the parameter configuration presented in the main text. R2 [min, max] = [0.48, 0.85]; observations [min, max] = [552761, 555880].
@@
Construction year: 1955-1960 (ref.: <<1955) 0.021 0.003 ***
Construction year: 1960-1965 (ref.: <<1955) -0.000 0.004
Construction year: 1965-1970 (ref.: <<1955) 0.045 0.003 ***
Construction year: 1970-1975 (ref.: <<1955) 0.050 0.004 ***
Construction year: 1975-1985 (ref.: <<1955) 0.084 0.004 ***
Construction year: 1985-1995 (ref.: <<1955) 0.175 0.004 ***
Construction year: 1995-2005 (ref.: <<1955) 0.270 0.004 ***
Construction year: 2005-2015 (ref.: <<1955) 0.343 0.004 ***
Construction year: 2015-2025 (ref.: <<1955) 0.404 0.006 ***
Construction year: Missing (ref.: <<1955) 0.046 0.013 ***
@@
To validate our definition of the historical memory of regions we repeated the hedonic regression analyses in Table 2 after considering an alternative dataset for past flood events that we compiled from ISPRA records.
EM-DAT is well-known to over-represent disasters in richer areas (Felbermayr and Gröschl, 2014). Therefore, we compiled a dataset of flood events from ISPRA data and re-estimated our hedonic regressions. ISPRA collects data on flood events drawing on multiple sources, such as the Italian Civil Protection Department, ISTAT, the National Research Council, local administrations and newspapers. For each event, it reports the duration, an estimation of the damages, an estimation of the allocated aid funds, and the total number of deaths and missing or evacuated people. Although data are provided back to 1951, systematic data collection only began in 2003, so we only consider the later period for better accuracy. Data are not collected in tabular format, but we manually compiled a database reading from the yearly reports by ISPRA. [^50] https://indicatoriambientali.isprambiente.it/it/edizioni-annuario. The latest available report is for the year 2022. The ISPRA dataset contains 402 event-region pairs distributed across regions and macroareas. Since the inclusion criteria for ISPRA are less strict than EM-DAT, this results on some variations in the historical memory classifications. Even with these differences we still observe the same pattern of significance of the risk coefficients disaggregated by awareness conditions that was observed in the regression presented in the main manuscript (Table A8).
Table A7: Flood risk effect on home prices disaggregated across terciles of awareness at the time of the transaction for different spatial fixed effects and half life times ($\tau$, see Section 2.4). $^{***}p<$ 0.01, $^{**}p<$ 0.05, and $^{*}p<$ 0.1. The awareness values are computed considering the flood events collected by (EM-DAT, 2024).
Dep. var.: log(P)
Model id 1 2 3 4 5 6 7 8 9
Half life
time ($\tau$)
7 10 17 7 10 17 7 10 17
Risk $\times$ High
awareness
-0.040∗∗
-0.038∗∗
-0.032∗∗
-0.016∗∗∗
-0.017∗∗∗
-0.014∗∗∗
-0.012∗∗∗
-0.012∗∗∗
-0.009∗∗
Risk $\times$ Medium
awareness
-0.029∗∗∗
-0.031∗∗∗
-0.036∗∗∗
-0.008
-0.006
-0.010∗∗∗
-0.007
-0.007
-0.010∗∗
Risk $\times$ Low
awareness
-0.013
-0.015
-0.019
-0.003
-0.002
-0.002
-0.005
-0.005
-0.006
Observations 552728 552728 552728
R2 0.669 0.755 0.848
Spatial fix. effects Municipality OMI zone Census Tract
N. clusters 7203 20297 160630
Clustered s.e.
Table A8: Flood risk effect on home prices. For each spatial fixed effects, Models 1-3-5 report the effect of the risk on home prices estimated using Equation (2), while Models 2-4-6 disaggregate the effect of flood risk by the level of flood awareness at the time of the purchase, Equation (4). Awareness is estimated following the procedure described in 2.4 using the data reported by ISPRA and $\tau=10$ years. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
Risk
-0.029∗∗∗
-0.010∗∗∗
-0.008∗∗
Risk $\times$ High
awareness
-0.038∗∗
-0.018∗∗∗
-0.014∗∗∗
Risk $\times$ Medium
awareness
-0.029∗∗∗
-0.005
-0.007
Risk $\times$ Low
awareness
-0.016
-0.002
-0.002
Observations 552728 552728 552728
@@
Spatial fix. effects Municipality OMI zone Census Tract
N. clusters 7203 20297 160630